### v1

# Exploring Training and Inference Scaling Laws in Generative Retrieval

Abstract.
Generative retrieval has emerged as a novel paradigm that leverages large language models (LLMs) to autoregressively generate document identifiers. Although promising, the mechanisms that underpin its performance and scalability remain largely unclear. We conduct a systematic investigation of training and inference scaling laws in generative retrieval, exploring how model size, training data scale, and inference-time compute jointly influence retrieval performance. To address the lack of suitable metrics, we propose a novel evaluation measure inspired by contrastive entropy and generation loss, providing a continuous performance signal that enables robust comparisons across diverse generative retrieval methods. Our experiments show that n-gram-based methods demonstrate strong alignment with both training and inference scaling laws, especially when paired with larger LLMs. Furthermore, increasing inference computation yields substantial performance gains, revealing that generative retrieval can significantly benefit from higher compute budgets at inference. Across these settings, LLaMA models consistently outperform T5 models, suggesting a particular advantage for larger decoder-only models in generative retrieval. Taken together, our findings underscore that model sizes, data availability, and inference computation interact to unlock the full potential of generative retrieval, offering new insights for designing and optimizing future systems. We release code at SLGR GitHub repository.
Generative Retrieval, Neural Scaling Law, Large Language Models
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Retrieval models and ranking
1. Introduction
Document retrieval is a fundamental area in information retrieval, focusing on retrieving relevant documents from large-scale corpora in response to user queries. Early retrieval systems were built on term-based heuristic methods, such as TF-IDF (Salton et al., 1975) and BM25 (Robertson and Zaragoza, 2009), which rely on query and document term overlap. With the development of pre-trained language models, like BERT (Devlin et al., 2019), retrieval evolved into the dense retrieval paradigm, where queries and documents are mapped into a shared high-dimensional vector space, achieving advanced performance in document retrieval. Recently, with the rise of generative large language models (LLMs) (OpenAI et al., 2024; Touvron et al., 2023; Raffel et al., 2020), a new paradigm called generative retrieval has emerged. Instead of matching queries with documents, generative retrieval directly generates documents based on a given query. By reformulating the retrieval task as an autoregressive generation problem, generative retrieval indeed introduces a novel solution to the research field.
Figure 1. Scaling laws of n-gram-based methods in training and inference. (1) Training Scaling Law: Contrastive Generation Loss shows a power-law relationship with model size for both T5 and LLaMA models. (2) Inference Scaling Law: Miss Rate decreases consistently with increasing inference FLOPs, demonstrating the scalability of n-gram-based methods across the T5-Base and LLaMA-7B model.
A central challenge in generative retrieval lies in designing effective document identifiers to represent documents, as generating entire long documents is impractical. The current identifiers can be divided into two broad categories based on how they carry semantics. 1) Natural identifiers retain inherent semantic information by leveraging components like titles (De Cao et al., 2021; Lee et al., 2022b; Chen et al., 2022; Li et al., 2023a) or n-gram fragments (Bevilacqua et al., 2022; Chen et al., 2023; Li et al., 2023b; Wang et al., 2023) extracted from the original text. Titles provide concise, human-readable overviews, while n-gram snippets capture more granular semantic features. 2) Learned identifiers, on the other hand, derive semantic representations through clustering or codebook methods. Notable examples include numeric IDs (Tay et al., 2022; Wang et al., 2022; Zhuang et al., 2023a; Mehta et al., 2023) and codebook-derived tokens (Yang et al., 2023; Sun et al., 2023; Zeng et al., 2024; Zhang et al., 2024c), which discretize document embeddings into token sequences. Existing generative retrieval methods within these two categories have continued to evolve, showing promising performance.
Despite these advancements, the core advantages of generative retrieval remain unclear, with no established consensus in the research community. One key reason for success in many LLM-based tasks is scaling—increasing model size, data volume, and inference computation (Brown et al., 2020; Wei et al., 2022; Wu et al., 2024). Given that generative retrieval follows the same autoregressive paradigm and is even built on LLM backbones, it is much more meaningful to explore scaling laws in generative retrieval to unlock the full potential of this paradigm.
While scaling laws have been extensively studied in various domains (Zhai et al., 2022; Radford et al., 2023; Jia et al., 2021; Radford et al., 2021; Zhai et al., 2025; Zhang et al., 2024b), exploring scaling in generative retrieval remains highly challenging. 1) To date, most studies have used relatively small encoder-decoder architectures (e.g., BART (Lewis et al., 2019), T5 (Raffel et al., 2020)) rather than larger, modern LLMs like LLaMA (Touvron et al., 2023). 2) Moreover, recent breakthroughs in LLM scaling have centered on decoder-only models—highly effective for generative tasks yet rarely explored for retrieval. 3) Standard retrieval metrics (e.g., recall, NDCG) are discrete and may miss nuanced performance variations, while contrastive entropy metrics from dense retrieval (Fang et al., 2024) are not suited for generative setups lacking a direct query-document scoring mechanism. 4) Additional complexity arises from the diverse ways of constructing document identifiers—whether natural or learned identifiers—each of which may respond differently to model scaling.
To address the above challenges, we introduce a new metric and employ larger models across different retrieval methods to systematically investigate how model size, training data scale, and inference compute impact performance. To capture subtle performance variations beyond discrete metrics, we propose a novel evaluation metric inspired by contrastive entropy and generation loss in neural scaling laws (Kaplan et al., 2020). Our metric measures the probability of generating the correct document identifier for a given query while considering random negative samples, yielding a continuous and sensitive performance signal. This metric enables consistent comparisons across various generative retrieval methods, models, and data scales. Leveraging this metric, we conduct extensive experiments to uncover the scaling behaviors of generative retrieval under different model sizes and data sizes. Additionally, we analyze how increased inference-time computation influences recall, providing further insight into the trade-offs and advantages of scaling computation.
From our extensive exploration, several intriguing findings stand out. 1) We observe that n-gram-based generative retrieval aligns remarkably well with both training and inference scaling laws, as illustrated in Figure 1, which presents clear scaling curves under varying model sizes and inference computation. 2) Expanding the training data scale benefits all methods, and n-gram-based approaches demonstrate especially robust gains, indicating a strong synergy between LLMs and natural identifiers. 3) We find that boosting inference computation (e.g., FLOPs) yields clear performance improvements that follow power-law trends, revealing that generative retrieval can significantly profit from additional inference computation—an aspect rarely discussed in prior work. 4) We discover that LLaMA models consistently outperform T5 models and exhibit higher theoretical upper bounds, hinting that the generative ability of larger decoder-only models may be particularly advantageous for generative retrieval.
2. Related Work
In this section, we revisit the background of generative retrieval and scaling laws. We first present the key advancements in generative retrieval, focusing on document identifier design and training methodologies. Then, we discuss neural scaling laws and their applications in various domains, including retrieval tasks.
2.1. Generative Retrieval
Generative retrieval leverages generative models, particularly LLMs, reformulating the retrieval task as an auto-regressive generation problem. Document identifiers, as the core component, should be distinctive enough to represent related passages, reducing useless information in a passage and making it easier for the model to memorize and learn. The design of DocID can be broadly categorized into two approaches: natural identifiers and learned identifiers.
The first type uses identifiers that naturally carry semantic meaning about its associated document, which are comparably cost-efficient to be created, without the need for additional human supervision or forcing any structure in the search space. A line of work (De Cao et al., 2021; Lee et al., 2022b; Chen et al., 2022; Li et al., 2023a) explored titles as identifiers and DSI(Tay et al., 2022) tested the first N words of the passage, providing human-readable summaries of document content. Multiple textual fragments  (Bevilacqua et al., 2022; Chen et al., 2023; Li et al., 2023b; Wang et al., 2023) extracted directly from the document are then proved to capture richer semantic features of the document. These n-grams are not tied to a specific document but reflect shared semantic content across multiple semantically related documents. The second type called learned identifiers, acquires semantic meaning through dense representation of passages clustering or codebook training, which induces structure in the search space with semantically similar documents having more similar document IDs. In this case, each document is assigned a unique numerical representation, and retrieval is achieved by generating the corresponding identifier. Specifically, numeric ID (Tay et al., 2022; Wang et al., 2022; Zhuang et al., 2023a; Mehta et al., 2023) was found easy to construct but required extra memory steps. While codebook-derived tokens (Yang et al., 2023; Sun et al., 2023; Zeng et al., 2024; Zhang et al., 2024c) have been proven effective because the search space is reduced after each decoding step.
Generative retrieval research has also advanced in training methods, which can be divided into generative training and discriminative training. In generative training, the model is trained to generate the appropriate identifier for a given query (Pradeep et al., 2023; Tay et al., 2022; Wang et al., 2022; Zeng et al., 2024; Bevilacqua et al., 2022), aligning naturally with the generative capabilities of LLMs to produce accurate document identifiers. Discriminative training, on the other hand, employs ranking losses (Li et al., 2024; Tang et al., 2024) and negative sample mining techniques (Zeng et al., 2024) to teach the model to produce a ranked list of documents. These approaches directly optimize retrieval performance by aligning the training objectives with the ranking requirements of retrieval tasks.
2.2. Neural Scaling Laws
Neural scaling laws describe predictable patterns of performance improvement as model size, dataset size, and computational resources increase. Baidu (Hestness et al., 2017) first introduced power-law relationships between test loss and these factors, offering an insight to predict neural network training. OpenAI (Kaplan et al., 2020) extended this concept to larger models, demonstrating that model scaling yields consistent improvements in tasks like language modeling. Google (Hoffmann et al., 2022) further advanced this understanding by introducing a unified formula for scaling laws, thus laying the groundwork for scaling strategies in neural networks.
Scaling laws have been successfully applied to various domain-specific fields such as speech recognition (Radford et al., 2023), computer vision (Dehghani et al., 2023; Zhai et al., 2022), and vision-language models (Jia et al., 2021; Radford et al., 2021). In the field of Information Retrieval, scaling laws have been explored in recommendation. Studies have examined applications in Click-Through Rate (CTR) prediction (Ardalani et al., 2022) and sequential recommendation models (Zhang et al., 2024a) using unique item identifiers. Recent research has demonstrated the effectiveness of trillion-parameter sequential transducers for generative recommendations (Zhai et al., 2025) and the development of architectures like Wukong (Zhang et al., 2024b) has established scaling laws for large-scale recommendation systems by effectively capturing diverse, high-order interactions through scalable network layers.
However, research on scaling laws in retrieval tasks remains limited. Scaling laws for dense retrieval (Fang et al., 2024) have been investigated, focusing on embedding-based methods using BERT-like models (Devlin et al., 2019). Another study on industrial multi-stage advertisement retrieval systems (Wang et al., 2024) emphasizes task-specific optimizations but lacks exploration of general scaling laws applicable to generative retrieval. In contrast, an early exploration of generative retrieval examines performance across varying corpus and model sizes (Pradeep et al., 2023) but is limited to learned identifier methods, leaving natural identifier methods unaddressed. Additionally, it only adopts different sizes of T5 models and does not fit power-law scaling relationships. These gaps motivate our comprehensive study, which examines generative retrieval under varying model sizes, data scales, and DocID strategies, aiming to shed light on broader scaling behaviors and guide future system design.
Following these discussions on generative retrieval and neural scaling laws, in this work, we conduct a comprehensive investigation of scaling behaviors in generative retrieval. By including larger decoder-only models and analyzing representative approaches, we aim to uncover key principles of scaling and provide insights to advance the design and optimization of generative retrieval systems.
3. Methodology
In this section, we first formalize the generative retrieval task in Section 3.1. Next, we describe the representative approaches for natural identifier and learned identifier: MINDER (Li et al., 2023b) and RIPOR (Zeng et al., 2024), respectively, in Section 3.2. The backbones and training configurations used for generative retrieval are introduced in Section 3.3. Finally, to address the limitations of traditional ranking metrics in reflecting the scaling behaviors of generative retrieval, we propose contrastive generation loss in Section 3.4.
3.1. Problem Formulation
We formalize the generative retrieval task as a two-step process. 1) Identifier assignment: assigning identifiers to documents and 2) Identifier Generation: generating query-specific identifiers to retrieve relevant documents.
Identifier Assignment. Let a corpus of documents be denoted as $\mathcal{D}=\{d_{1},d_{2},\ldots,d_{N}\}$. Each document $d\in\mathcal{D}$ is associated with a set of identifiers $\mathcal{I}_{d}=\{i_{d}^{1},i_{d}^{2},\ldots,i_{d}^{M}\}$ through a transformation function $h(d;\psi)$, parameterized by $\psi$:
(1) $\mathcal{I}_{d}=h(d;\psi).$
Identifier Generation. During training, the generative language models $f$ parameterized by $\theta$, learn to generate query-specific identifiers $\mathcal{I}_{q}=\{i_{q}^{1},i_{q}^{2},\ldots,i_{q}^{K}\}$ based on paired training data consisting of queries and relevant document identifiers. During inference, given a query $q$, the model $f$ use $q$ as input and generate a set of identifiers:
(2) $\mathcal{I}_{q}=f(q;\theta).$
The generated identifiers $\mathcal{I}_{q}$ are then used to retrieve documents $\mathcal{D}_{q}$ by applying the reverse of the identifier assignment function.
Different generative approaches usually focus on designing different types of identifiers. However, the goal remains the same: to optimize the transformation function $h(d;\psi)$ and the generative language model $f(q;\theta)$ to generate query-specific identifiers that maximize the relevance of the retrieved documents $\mathcal{D}_{q}$.
3.2. Representative Generative Retrieval Methods
As outlined above, the document identifier is a crucial component of generative retrieval. There are two primary categories: natural identifiers and learned identifiers. In this study, we select representative techniques from each category to evaluate their scaling behaviors. For natural identifiers, we use the n-gram-based method as n-grams effectively capture diverse semantic relationships between queries and documents. For learned identifiers, we adopt a codebook-based approach, as it leverages advanced neural methods to encode semantic information more compactly and effectively.
3.2.1. N-gram-based Generative Retrieval
N-grams offer a flexible and query-adaptive approach to document identifiers (Bevilacqua et al., 2022; Chen et al., 2023; Li et al., 2023b; Wang et al., 2023). These identifiers are directly extracted from documents based on their overlap or semantic relevance to specific queries, capturing key contextual features that align closely with user queries.
In this method, an LLM is trained to generate query-specific n-grams, which act as identifiers for ranking documents. The training process begins by selecting n-grams from documents based on their overlap or semantic similarity with user queries. These n-grams serve as the basis for training the LLM to predict relevant identifiers when presented with a query. At inference, the LLM generates n-grams for a query, and these are used to score and rank documents based on a heuristic function, such as n-gram frequency or semantic similarity.
For our experiments, we adopt the method of MINDER (Li et al., 2023b), focusing on extracting identifiers from the body text of documents.
3.2.2. Codebook-Based Generative Retrieval
Codebooks originate from techniques designed to create discrete visual data representations (Lee et al., 2022a; van den Oord et al., 2017). Learned codebooks for documents represent them as sequences of unique codes that effectively capture the semantics of their associated content (Yang et al., 2023; Sun et al., 2023; Zeng et al., 2024; Zhang et al., 2024c).
The process begins by encoding documents into dense vector representations using an encoder network. These dense vectors are then discretized into tokens by mapping them to entries in the learned codebook. Finally, a decoder network reconstructs the original document from the codebook tokens to ensure the generated representations are accurate and compact. The resulting code sequences serve as unique identifiers for documents, establishing a one-to-one correspondence between the sequence and the document. At inference, the LLM generates a code sequence for a query, which is matched to the corresponding document based on the learned codebook.
In our study, we select the RIPOR (Zeng et al., 2024) as the representative of the codebook-based method.
3.3. Backbone Models and Training Setting
To investigate the scaling capabilities of different generative retrieval systems, it is essential to first identify the backbone models they utilize and their corresponding training settings.
3.3.1. Backbone Models
For our experiments, we use the widely adopted T5 (Raffel et al., 2020) series as the primary backbone, which has been extensively employed in previous generative retrieval studies (Tay et al., 2022; Zeng et al., 2024; Ren et al., 2023; Mehta et al., 2023; Zhuang et al., 2023b; Zhang et al., 2024c). To evaluate the effect of model size, we experiment with all T5 variants: T5-Small, T5-Base, T5-Large, T5-XL, and T5-XXL, which differ only in parameter sizes while maintaining identical pre-training configurations.
To further investigate architectures beyond encoder-decoder models and explore the impact of larger parameter scales, we also experiment with LLaMA models. Specifically, we employ three sizes of LLaMA-2 (Touvron et al., 2023) models: LLaMA-2-7B, LLaMA-2-13B, and LLaMA-2-70B. These decoder-only models not only represent a different architectural paradigm but also significantly increase the parameter scale compared to typical encoder-decoder backbones. Their scaling properties observed in other tasks provide valuable insights into how architecture and model size influence generative retrieval performance.
3.3.2. Training Setting
The training process involves two distinct setups corresponding to the two generative retrieval approaches: n-gram-based generative retrieval and codebook-based generative retrieval. In the following sections, we detail the datasets, training configurations, and loss functions for the two representative approaches, respectively.
$\bullet$ N-Gram-Based Generative Retrieval. For n-gram-based generative retrieval, we use the Natural Questions (NQ) (Kwiatkowski et al., 2019) dataset, which contains over 20 million documents. Following the MINDER methodology, which originally involves three types of identifiers, we focus exclusively on the body text as the identifier type. This simplification is made because our study emphasizes scaling behavior rather than achieving absolute performance, and the body text serves as the most important identifier type. For each document, we select 10 n-grams, each consisting of 10 tokens, based on their overlap with the associated query to ensure high semantic relevance. The final training set consists of nearly 600,000 query-to-n-gram pairs.
During training, the input consists of the query, and the label is a single n-gram from the corresponding document. The training objective is to minimize the cross-entropy loss for generating each n-gram $n$, given the query $q$:
(3) $\mathcal{L}_{\text{n-gram}}=-\log P(n\mid q;\theta),$
where $n$ represents an individual n-gram and $\theta$ represents the model parameters. Since each n-gram is treated independently, the model learns to predict each query-relevant n-gram as a separate target during training.
To keep consistent with the following setup of codebook-based methods, we train each model for one epoch using the MINDER-provided data with LoRA (Hu et al., 2022). As recommended in MINDER, the learning rate for T5 models is set to 3e-5. For LLaMA models, we use a learning rate of 3e-4, a commonly adopted value for fine-tuning LLaMA.
$\bullet$ Codebook-Based Generative Retrieval. Following the selected representative codebook-based generative retrieval method, RIPOR (Zeng et al., 2024), we conduct experiments using the MSMARCO-1M dataset, a subset of the MSMARCO (Nguyen et al., 2017) dataset, containing one million passages and query-document pairs. This dataset is chosen because RIPOR is only available on MSMARCO, and since our study focuses on scaling behavior rather than absolute performance, the choice of the dataset has minimal impact on the relative results.
The codebook consists of $N_{c}$ unique codes, with each document represented as a sequence of $L_{c}$ codes. These codes are treated as new tokens added to the vocabulary of the LLM $\theta$. The training follows a standard generative setup, where the input is the query and the output is the document’s code sequence. The training objective is to minimize the cross-entropy loss, which measures the negative log-likelihood of generating the correct code sequence $\mathbf{c}_{d}=\{c_{1},c_{2},\ldots,c_{L_{c}}\}$ for document $d$, given query $q$:
(4) $\mathcal{L}_{\text{codebook}}=-\sum_{t=1}^{L_{c}}\log P(c_{t}\mid q,c_{<t};\theta).$
In our experiments, $N_{c}=256$ unique codes, and $L_{c}=32$ codes per document. We train each model for one epoch using the RIPOR-provided data with LoRA (Hu et al., 2022). For T5 models, the learning rate is set to 1e-3, consistent with configurations in RIPOR. For LLaMA models, the learning rate is set to 3e-4.
3.4. Evaluation
Evaluating generative retrieval models requires metrics that effectively capture nuanced variations of retrieval performance. Traditional retrieval metrics, such as NDCG and Recall, are not well-suited for this purpose. First, these metrics are inherently discrete, making them incapable of capturing fine-grained differences in model outputs. Second, they primarily evaluate changes in ranked lists, offering limited insight into the nuanced behavior of model outputs. Lastly, their reliance on cutoff parameters means they only consider documents within a specific range (e.g., top $K$), ignoring contributions from documents ranked lower, which limits their effectiveness for studying scaling behaviors in generative retrieval.
To address these limitations, we draw inspiration from prior work on dense retrieval metrics (Fang et al., 2024) and scaling laws in large language models (Kaplan et al., 2020) to propose a novel evaluation metric tailored for generative retrieval. Building on the concept of contrastive entropy used in dense retrieval, we adapt it to the generative setting by incorporating the loss associated with query-to-identifier generation.
In generative retrieval, the primary objective is to generate identifiers that correspond to relevant documents while distinguishing them from irrelevant ones. To quantify this ability, we define a contrastive generation loss (CGL), which evaluates the model’s capacity to generate identifiers for positive documents in the presence of negative documents.
For a query $q$ and its associated positive document $d^{+}$, let $\mathcal{I}_{d^{+}}$ be the identifier for $d^{+}$, and let $\mathcal{I}_{d^{-}}$ be the identifiers for a set of negative documents $\mathcal{D}^{-}$. The contrastive generation loss $\mathcal{L_{\text{CGL}}}$ is defined as:
(5) $\mathcal{L}_{\text{CGL}}=-\log\frac{\sum_{d^{-}}\mathcal{L}(q,\mathcal{I}_{d^{-}})}{\mathcal{L}(q,\mathcal{I}_{d^{+}})+\sum_{d^{-}}\mathcal{L}(q,\mathcal{I}_{d^{-}})},$
where $\mathcal{L}(q,\mathcal{I})$ represents the generation loss for the query $q$ to produce the identifier $\mathcal{I}$, calculated using the cross-entropy loss.
The contrastive generation loss is a continuous and highly sensitive metric designed to effectively evaluate retrieval performance across different generative retrieval methods and models. 1) For n-gram-based generative retrieval, where each document is represented by multiple identifiers, the $\mathcal{L}(q,\mathcal{I}_{d^{+}})$ is extended by averaging the generation loss across all n-grams associated with the positive document. 2) For codebook-based generative retrieval, the query-to-document relationship is evaluated through the loss of generating the document’s unique code sequence. 3) Furthermore, this metric is versatile across different LLM architectures, as it directly leverages the generative capabilities of the models to compute losses. 4) Moreover, the CGL is based on the relative proportion of the positive document’s loss compared to negative documents, reducing sensitivity to the absolute magnitude of the loss and ensuring consistency across varying model scales and training settings.
To validate the proposed CGL, we conducted experiments to analyze its relationship with existing ranking metrics, such as Recall, NDCG, MAP, and MRR. Using the MINDER framework as an example, we trained models of varying sizes and configurations and then calculated their retrieval performance using Recall@100, NDCG@10, MAP@10, MRR@10 as well as the proposed CGL. Results in Figure 2 reveal an almost linear relationship between contrastive generation loss and standard metrics. This alignment demonstrates that the proposed metric effectively captures retrieval performance, offering a consistent and reliable evaluation framework across different settings.
Figure 2. Relationship between proposed Contrastive Generation Loss and traditional retrieval metrics (Recall@100, NDCG@10, MRR@10, and MAP@10). The results demonstrate a nearly linear correlation, validating the effectiveness of CGL in reflecting retrieval performance as measured by traditional metrics.
Figure 3. Scaling behavior of contrastive generation loss regarding model size for (1) n-gram-based and (2) codebook-based method. The results demonstrate a clear scaling trend for the n-gram-based approach, while the codebook-based method exhibits no consistent improvement as model size increases.
4. Training Scaling Laws
In this section, we present the results of our experiments and summarize our initial investigation into the scaling laws for generative retrieval. Specifically, we analyze how model size, training data size, and identifier methods influence retrieval performance, using contrastive generation loss as the evaluation metric. These experiments aim to uncover the relationships and trends that define the scalability of generative retrieval systems.
4.1. Model Size Scaling
We now investigate the impact of model size on generative retrieval performance. We begin by examining n-gram-based generative retrieval.
4.1.1. N-Gram-Based Generative Retrieval
Table 1. Fitted parameters for the scaling law on model sizes with n-gram-based methods.
Method Model $\gamma$ $\alpha$ $\lambda_{P}$ $R^{2}$
N-Gram T5 Series $2.26\times 10^{-2}$ $0.40$ $0.00356$ $0.996$
N-Gram LLaMA Series $1.24\times 10^{8}$ $2.40$ $0.00328$ $0.999$
We investigated the effect of model size on generative retrieval performance using the n-gram-based method. Models of varying sizes, including T5 and LLaMA, were fine-tuned on query-to-n-gram training pairs, and their performance was evaluated using the contrastive generation loss on the test set. Figure 3 (1) illustrates the scaling behavior of T5 and LLaMA models concerning this metric. Based on the observed relationship between model size and the contrastive generation loss, we propose a scaling law to quantify this behavior as follows:
(6) $\mathcal{L}_{\text{CGL}}(P)=\left(\frac{\gamma}{P}\right)^{\alpha}+\lambda_{P}.$
Here, $P$ represents the number of non-embedding parameters of the model, and $\mathcal{L}_{\text{CGL}}(P)$ denotes the contrastive generation loss on the test set. The parameters $\gamma$, $\alpha$, and $\lambda_{P}$ are coefficients determined through fitting. Here, $\lambda_{P}$ represents the irreducible loss, a theoretical lower bound on performance as $P$ approaches infinity, accounting for limitations such as dataset noise and variability in relevance judgments.
Using the least squares method, we fit the scaling law and report the fitted coefficients for T5 and LLaMA in Table 1. The results reveal several important insights: 1) Both models demonstrate a strong power-law relationship between model size and contrastive generation loss, with exceptionally high coefficients of determination (T5: R² = 0.996, LLaMA: R² = 0.999). 2) LLaMA demonstrates comprehensive performance advantages over T5, characterized by a more efficient scaling mechanism. Specifically, LLaMA achieves lower contrastive generation loss across model sizes and exhibits a steeper improvement curve (scaling exponent of $\alpha=2.40$ versus T5’s $\alpha=0.40$). Notably, LLaMA’s smaller irreducible loss ($\lambda_{P}$) suggests a higher potential performance ceiling, indicating its superior capability to approach the theoretical limits of generative retrieval performance. These findings highlight LLaMA’s promising performance and potentially signal the broader potential of decoder-only architectures.
4.1.2. Codebook-Based Generative Retrieval
For the codebook-based method, we conducted similar experiments, fine-tuning different sizes of T5 and LLaMA models on query-to-code sequence training pairs and evaluating their performance using contrastive generation loss on the test set. The results, shown in Figure 3 (2), reveal that neither T5 nor LLaMA models exhibit a consistent reduction in CGL as the model size increases, with the value fluctuating across different model sizes and showing no clear scaling trend. This suggests that increasing model size does not inherently enhance retrieval performance for codebook-based methods. This finding aligns with prior research (Pradeep et al., 2023), where T5-XXL underperformed smaller T5-XL with similar generative retrieval methods.
The possible reasons are as follows: 1) Codebook tokens are newly introduced and unrelated to the models’ pretraining objectives, requiring the models to learn entirely new semantic relationships during fine-tuning. 2) Newly introduced tokens often demand more extensive training to be fully integrated into the model’s generative capabilities. In our experiments, only a single epoch of fine-tuning was conducted, which may not have been sufficient for the models to fully learn the codebook representations. Scaling behavior might emerge with additional training epochs once the models better understand these novel tokens. We leave this possibility for future research as the computational intensity needed makes comprehensive exploration impractical for us.
Despite the lack of scaling trends, LLaMA consistently achieves lower contrastive generation loss than T5 across all model sizes, highlighting its stronger retrieval capabilities. These findings suggest LLaMA demonstrates promising performance characteristics and may signal avenues for future research in generative retrieval.
Figure 4. Scaling behavior of contrastive generation loss concerning training data size for n-gram-based (left) and codebook-based (right) methods using the LLaMA-7B model. The results show clear scaling trends for both methods, with a steeper decline observed for n-grams-based generative retrieval.
4.1.3. Comparsions.
The results presented in Figure 3 highlight clear differences between the n-gram-based and codebook-based methods in terms of scaling behavior and overall performance. 1) The n-gram-based method significantly outperforms the codebook-based approach, demonstrating substantially lower contrastive generation loss. Even the most advanced LLaMA models utilizing codebook tokens cannot match the performance of T5 models with n-gram-based retrieval. This performance gap highlights the intrinsic challenges of codebook tokens, which lack the semantic coherence and natural language alignment inherent in n-grams. 2) LLaMA consistently outperforms T5 across both methods, achieving lower CGL at comparable model sizes. This highlights LLaMA’s stronger generative capabilities and its architectural advantage.
4.2. Data Size Scaling
The size of the training dataset also plays a critical role in determining the performance of generative retrieval models. In this section, we investigate how varying the training data size influences retrieval performance while keeping other factors, such as model size and architecture, constant.
Table 2. Fitted parameters for the scaling law on data sizes.
Method Model $\eta$ $\beta$ $\lambda_{D}$ $R^{2}$
N-Gram LLaMA-7B $1.05\times 10^{4}$ $3.99$ $0.00335$ $0.990$
Codebook LLaMA-7B $9.28\times 10^{-1}$ $1.01$ $0.00389$ $0.983$
4.2.1. N-Gram-Based Generative Retrieval
To study the effect of training data size, we use the LLaMA-2-7B model and incrementally increase the number of training pairs constructed using the n-gram-based method. Figure 4 shows the scaling behavior of the contrastive generation loss ($\mathcal{L}_{\text{CGL}}$) concerning training data size. Similar to model size, we fit the scaling behavior using the following power-law equation:
(7) $\mathcal{L}_{\text{CGL}}(D)=\left(\frac{\eta}{D}\right)^{\beta}+\lambda_{D}$
Here, $D$ represents the number of query-identifier pairs, and $\mathcal{L}_{\text{CGL}}(D)$ denotes the contrastive generation loss on the test set. The parameters $\eta$, $\beta$, and $\lambda_{D}$ are coefficients determined through fitting. The term $\lambda_{D}$ represents the irreducible loss, a theoretical lower bound on retrieval performance as $D$ approaches infinity.
Using the least squares method, we fit the scaling law to the observed data, achieving a coefficient of determination of $R^{2}=0.990$, which indicates a strong fit. As seen in Figure 4 (1), retrieval performance improves significantly as the training data size increases, with the contrastive generation loss decreasing sharply. The power-law scaling behavior reflects the model’s capacity to leverage larger datasets to better capture the semantic relationships between queries and identifiers.
4.2.2. Codebook-Based Generative Retrieval
We also evaluated the effect of training data size on retrieval performance for the codebook-based method. Similar to the n-gram-based experiment, the LLaMA-2-7B model was fine-tuned on training datasets of varying sizes, and constructed with query-code sequence pairs.
Using the same power-law equation as for n-grams, we fit the scaling behavior of the codebook-based method. The fitted curve in Figure 4 (2) achieves a coefficient of determination of $R^{2}=0.983$, indicating a strong fit. As the training data size increases, the contrastive generation loss decreases steadily, demonstrating that retrieval performance improves with larger datasets. The results highlight that even for the codebook-based method, which involves learning entirely new representations unrelated to the model’s pretraining objectives, increasing the data size leads to significant retrieval performance enhancements.
4.2.3. Comparsions.
The results in Figure 4 highlight key differences between n-gram-based and codebook-based methods in their scaling behavior and overall retrieval performance.
For n-gram-based methods, the scaling exponent ($\beta=3.99$) is much larger than that of codebook-based methods ($\beta=1.01$), indicating a steeper improvement in performance with increased data size. This can be attributed to the semantic richness of n-grams, which align closely with the model’s pretraining objectives, allowing the model to fully leverage larger datasets. In contrast, the codebook-based method lacks such alignment, resulting in a slower rate of improvement as data size increases. The low scaling exponent implies that this method requires substantially more training data to achieve comparable performance. Recent studies suggest more advanced training strategies, such as ranking losses (Tang et al., 2024; Zeng et al., 2024), could potentially address these learning challenges.
4.3. Model-Data Joint Laws
To capture the joint effects of model size and data size on retrieval performance, we combine the observations from the previous sections into a single scaling function. Inspired by established scaling laws in LLMs (Kaplan et al., 2020), we employ the following equation to describe the combined effects:
(8) $\mathcal{L}(P,D)=\left(\left(\frac{\gamma}{P}\right)^{\frac{\alpha}{\beta}}+\frac{\eta}{D}\right)^{\beta}+\delta.$
Here, $P$ and $D$ represent the model size (number of non-embedding parameters) and training data size, respectively. The parameters $\gamma$, $\eta$, $\alpha$, $\beta$, and $\delta$ are coefficients determined through fitting. Based on experimental results using LLaMA with n-gram-based method across various model sizes and training data sizes, we obtained the following estimates for these coefficients:
(9) $\gamma=6.32\times 10^{3},\quad\alpha=3.27,\quad\beta=0.95,$
(10) $\eta=3.37\times 10^{5},\quad\delta=3.26\times 10^{-3},\quad R^{2}=0.976.$
The coefficient of determination indicates a high degree of accuracy in capturing the relationship between model size, data size, and retrieval performance. This unified scaling function highlights the complementary contributions of model size and data size to retrieval performance. Larger models reduce loss by better capturing semantic relationships, while increased data size allows for improved learning of these relationships. The joint law provides a valuable framework for balancing model size and data requirements to optimize performance efficiently.
5. Inference Scaling Laws
In the previous sections, we demonstrated the existence of scaling laws in generative retrieval concerning model size and data size, highlighting how larger models and richer data yield stronger performance. These findings align with trends observed in LLMs, further validating the synergy between generative retrieval and LLMs.
Recent studies have also revealed another dimension of scaling: the scaling of computational investment during inference (Wu et al., 2024). Specifically, increasing inference computing, such as scaling the number of decoding tokens, has been shown to yield substantial performance gains. This trend suggests that inference scaling may be a critical factor in optimizing LLM-based systems, complementing improvements from larger models and data.
In generative retrieval, inference scaling is particularly promising because the identifiers used for retrieval are generated dynamically at this stage. Generative retrieval inherently involves generating query-specific identifiers to score and rank documents. The quality and diversity of generated identifiers are directly influenced by the inference process, including parameters such as beam size, which controls the breadth of the search during generation. Larger beam sizes allow for the exploration of a broader search space, potentially improving retrieval performance, but they also increase inference latency, creating a trade-off between quality and efficiency.
Motivated by this, we investigate whether the inference scaling laws observed in LLMs extend to generative retrieval. Given that generative retrieval shares a strong alignment with the LLMs, we wonder whether similar scaling behaviors will hold for generative retrieval inference.
Figure 5. Inference scaling behavior of n-gram-based methods across T5 and LLaMA models. Miss Rate consistently decreases as inference FLOPs per query increase, demonstrating a power-law relationship. LLaMA models show a steeper decline, particularly at higher inference FLOPs, highlighting their superior scalability compared to T5.
5.1. Experimental Setup
To explore this, we focus on n-gram-based generative retrieval as a representative approach. N-gram-based methods are particularly suitable for investigating inference scaling because their retrieval process aligns well with the core principles of scaling during inference. Specifically, these methods first generate n-grams and then use these n-grams to score and rank documents. By increasing the beam size during n-gram generation, we can systematically expand the set of candidate n-grams. This increase not only boosts the quantity of n-grams available for scoring but also enhances their diversity and quality, which is likely to improve the final document retrieval performance.
We use the n-gram-based generative retrieval method described in Section 3.2.1. And we chose T5-Base and LLaMA-7B as representative models, both are fine-tuned under identical settings on the NQ dataset (Kwiatkowski et al., 2019), using the same query-to-n-gram pairs. Each model is trained for one epoch to ensure consistency across experiments.
During inference, we vary the beam size, testing values of $B=\{1,5,10,20,50,100\}$ and record the corresponding inference computation needed. Beam size determines the number and search space of candidate n-grams generated per query, which in turn affects the diversity of identifiers used for document retrieval. Increasing the beam size effectively increases the computational cost during inference, as a larger search space requires more floating-point operations (FLOPs) to generate candidate n-grams. To ensure comparability, the number of retrieved documents is fixed at 100 for all configurations, and other settings like batch size remain constant.
To evaluate retrieval performance, we define a metric called Miss Rate (MR), which measures the proportion of relevant documents that are not retrieved within the top $k$ results.
(11) $\text{MR}@k=1-\text{Recall}@k,$
where $k$ represents the number of retrieved documents considered (e.g., $k=5,20,100$). MR provides a straightforward and interpretable view of retrieval effectiveness by focusing on the proportion of relevant documents missed. By analyzing MR@5, MR@20, and MR@100, we systematically evaluate how varying beam sizes influence retrieval performance across different levels of precision.
Table 3. Fitted parameters for the inference scaling law with n-gram-based methods.
Model Miss Rate $\mu$ $\sigma$ $\lambda_{C}$ $R^{2}$
T5-Base MR@5 $1.60\times 10^{-4}$ $0.0620$ $0.3834$ $0.915$
T5-Base MR@20 $3.85\times 10^{-4}$ $0.0508$ $0.1276$ $0.970$
T5-Base MR@100 $1.71\times 10^{-2}$ $0.0755$ $0.0665$ $0.983$
LLaMA-7B MR@5 $2.71\times 10^{9}$ $0.3479$ $0.2779$ $0.999$
LLaMA-7B MR@20 $4.16\times 10^{9}$ $0.4233$ $0.1859$ $0.999$
LLaMA-7B MR@100 $4.90\times 10^{9}$ $0.4862$ $0.1141$ $0.999$
5.2. Results
The results of our experiments are summarized in Figure 5, which illustrates the Miss Rate across different inference FLOPs per query for both T5-Base and LLaMA-7B models. We evaluate MR at different retrieval thresholds ($k=5,20,100$) to assess the retrieval performance under varying levels of precision.
To analyze the relationship between inference computational cost and retrieval performance, we propose a fitting function:
(12) $\text{MR}(C)=\left(\frac{\mu}{C}\right)^{\sigma}+\lambda_{C},$
where $C$ represents the inference FLOPs per query, $\mu$, $\sigma$, and $\lambda_{C}$ are parameters to fit, and $\lambda_{C}$ is irreducible loss, a theoretical lower bound on retrieval performance as $C$ approaches infinity.
The fitted curves in Figure 5, reveal a consistent trend for both T5-Base and LLaMA-7B models: the Miss Rate decreases steadily as the inference FLOPs per query increase. This validates the effectiveness of inference scaling in improving generative retrieval performance. The decrease in MR is particularly pronounced at lower computational budgets, where increased beam sizes lead to significant performance gains.
As shown in Table 3, the proposed scaling law provides an excellent fit to the experimental data for both models, with high $R^{2}$ values in most cases. Additionally, the rate of decline in MR varies with the retrieval threshold $k$. For both models, MR@5 decreases the fastest, reflecting the substantial benefits of inference scaling for high-precision retrieval. Conversely, MR@100 exhibits a slower decline, indicating that broader recall settings are less sensitive to increased computational resources.
Comparing the two models, when inference FLOPs are below $10^{11}$, T5-Base outperforms LLaMA-7B, achieving lower Miss Rates across all $k$ values. However, as FLOPs increase beyond $10^{11}$, LLaMA-7B demonstrates a clear advantage, with all three Miss Rates achieving lower values. This highlights that LLaMA-7B gains substantially more performance rise from increased inference computation compared to T5-Base, making it highly effective in high-computation scenarios. Notably, LLaMA-7B’s irreducible loss ($\lambda_{C}$) for MR@5 is significantly lower than T5-Base, indicating superior potential for high-precision retrieval, while its irreducible loss for MR@20 and MR@100 is slightly higher, suggesting T5-Base retains an advantage in broader recall settings. These results highlight the trade-offs between the two models, with LLaMA-7B excelling in high-precision tasks, achieving lower Miss Rates for the top-ranked results as computational budgets grow, while T5-Base demonstrates greater potential for broader recall tasks like larger-scale retrieval.
6. Discussion
In this work, we explored the training scaling laws of model size, data size, and inference scaling laws in generative retrieval, proposing metrics to analyze retrieval performance under varying computational budgets. By systematically varying these factors, we highlighted how scaling influences retrieval performance and identified key trade-offs between inference computation and performance. While our findings provide valuable insights into generative retrieval systems, several important aspects remain unexplored and warrant further investigation.
We introduced a new metric contrastive generation loss to evaluate the generative retrieval capability of models. While CGL provides a relative measure of a model’s retrieval ability under the same settings, it is not equivalent to traditional metrics such as Recall. For example, two models with similar CGL values may exhibit differing recall performance, and a lower CGL does not guarantee higher recall. CGL is useful for comparing models in relative terms within the same experimental setup, but its interpretability as an absolute measure of retrieval performance is limited.
Another limitation of this work lies in the training methodology, as we relied solely on the basic generative cross-entropy loss. While this aligns well with n-gram-based methods, which match the model’s pretraining objective and require minimal fine-tuning, it poses challenges for codebook-based methods that need to learn entirely new tokens or relationships, demanding more training data and epochs. We attempted to use RIPOR’s full dataset, containing over 80 million query-to-code sequences from MSMARCO, but this scale of data was computationally infeasible for LLaMA, as it would require thousands of hours for us to complete the training. Future research with greater resources could further explore the potential of such extensive datasets and more advanced training objectives.
In our experiments, we observed that codebook-based methods did not exhibit a clear scaling effect with increasing model size, but showed a pronounced scaling effect with training data size. This discrepancy may stem from the significantly higher learning difficulty of codebook representations compared to n-grams. The amount of training data used in our experiments may have been insufficient, leaving the models in an underfitted state and preventing the scaling effect with model size from emerging. Previous studies have shown that model performance can undergo substantial improvement at a certain point (Du et al., 2024), it is possible that with sufficient training data and time, codebook-based methods might also exhibit scaling with model size. Future research with enhanced settings and resources could further investigate this possibility.
In this study, we primarily focused on the n-gram-based method for inference scaling. We observed improved retrieval performance by increasing beam size and validated its alignment with power-law scaling. However, codebook-based methods could also potentially benefit from inference scaling strategies, such as increasing beam size to explore a larger search space. Due to space limitations, we did not explore this direction, but it represents a possible direction for future research.
7. Conclusion and Future Work
In this paper, we explored the scaling laws of generative retrieval, analyzing how model size, training data size, model architecture, and retrieval methods impact retrieval performance. Using Contrastive Generation Loss as our primary metric, we found a clear power-law relationship between model size and retrieval performance for n-gram-based methods. Additionally, we observed consistent scaling trends with increasing data size for both n-gram-based and codebook-based methods. A comparison between architectures showed that LLaMA consistently outperformed T5 under identical experimental conditions, highlighting LLaMA’s superior capability for advancing generative retrieval.
Beyond model and data scaling, we also investigated inference scaling, specifically for n-gram-based methods. Our results revealed that increasing inference computation followed a distinct power-law trend, significantly boosting retrieval performance. This unique inference scaling offers complementary opportunities for optimizing retrieval efficiency and performance. Collectively, our findings validate the effectiveness of systematically scaling model capacity, training data, and inference computation, suggesting promising pathways to further enhance generative retrieval methods.
While our research provides a foundation for understanding the scaling laws of generative retrieval, several directions for future exploration remain. First, further investigation into scaling laws for codebook-based methods is warranted, particularly in the context of model size. Exploring advanced training objectives could better capture the potential scaling benefits of these methods, which were not fully realized under our experimental settings. Second, with access to greater computational resources, revisiting extensive datasets could enable deeper investigations into the scaling laws of generative retrieval, particularly for methods that require larger-scale training. Finally, while this study focused on n-gram-based methods for inference scaling, exploring similar scaling strategies for codebook-based methods, such as increasing beam size, remains an intriguing direction for future research.

### v2

@@
 Abstract.
-Generative retrieval has emerged as a novel paradigm that leverages large language models (LLMs) to autoregressively generate document identifiers.
-Although promising, the mechanisms that underpin its performance and scalability remain largely unclear.
-We conduct a systematic investigation of training and inference scaling laws in generative retrieval, exploring how model size, training data scale, and inference-time compute jointly influence retrieval performance.
-To address the lack of suitable metrics, we propose a novel evaluation measure inspired by contrastive entropy and generation loss, providing a continuous performance signal that enables robust comparisons across diverse generative retrieval methods.
-Our experiments show that n-gram-based methods demonstrate strong alignment with both training and inference scaling laws, especially when paired with larger LLMs.
-Furthermore, increasing inference computation yields substantial performance gains, revealing that generative retrieval can significantly benefit from higher compute budgets at inference.
+Generative retrieval reformulates retrieval as an autoregressive generation task, where large language models (LLMs) generate target documents directly from a query.
+As a novel paradigm, the mechanisms that underpin its performance and scalability remain largely unexplored.
+We systematically investigate training and inference scaling laws in generative retrieval, exploring how model size, training data scale, and inference-time compute jointly influence performance.
+We propose a novel evaluation metric inspired by contrastive entropy and generation loss, providing a continuous performance signal that enables robust comparisons across diverse generative retrieval methods.
+Our experiments show that n-gram-based methods align strongly with training and inference scaling laws.
+We find that increasing model size, training data scale, and inference-time compute all contribute to improved performance, highlighting the complementary roles of these factors in enhancing generative retrieval.
 Across these settings, LLaMA models consistently outperform T5 models, suggesting a particular advantage for larger decoder-only models in generative retrieval.
-Taken together, our findings underscore that model sizes, data availability, and inference computation interact to unlock the full potential of generative retrieval, offering new insights for designing and optimizing future systems.
+Our findings underscore that model sizes, data availability, and inference computation interact to unlock the full potential of generative retrieval, offering new insights for designing and optimizing future systems.
 We release code at SLGR GitHub repository.
 Generative Retrieval, Neural Scaling Law, Large Language Models
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Retrieval models and ranking
+journalyear: 2025copyright: ccconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3729973isbn: 979-8-4007-1592-1/2025/07ccs: Information systems Retrieval models and ranking
 1. Introduction
@@
 With the development of pre-trained language models, like BERT (Devlin et al., 2019), retrieval evolved into the dense retrieval paradigm, where queries and documents are mapped into a shared high-dimensional vector space, achieving advanced performance in document retrieval.
-Recently, with the rise of generative large language models (LLMs) (OpenAI et al., 2024; Touvron et al., 2023; Raffel et al., 2020), a new paradigm called generative retrieval has emerged.
+Recently, with the rise of generative large language models (LLMs) (OpenAI et al., 2024; Touvron et al., 2023; Raffel et al., 2020), a new paradigm called generative retrieval has emerged (Li et al., 2024a).
 Instead of matching queries with documents, generative retrieval directly generates documents based on a given query.
@@
 (1) Training Scaling Law: Contrastive Generation Loss shows a power-law relationship with model size for both T5 and LLaMA models.
-(2) Inference Scaling Law: Miss Rate decreases consistently with increasing inference FLOPs, demonstrating the scalability of n-gram-based methods across the T5-Base and LLaMA-7B model.
+(2) Inference Scaling Law: Miss Rate decreases consistently with increasing inference FLOPs across the T5-Base and LLaMA-7B model.
 A central challenge in generative retrieval lies in designing effective document identifiers to represent documents, as generating entire long documents is impractical.
@@
 2) Learned identifiers, on the other hand, derive semantic representations through clustering or codebook methods.
-Notable examples include numeric IDs (Tay et al., 2022; Wang et al., 2022; Zhuang et al., 2023a; Mehta et al., 2023) and codebook-derived tokens (Yang et al., 2023; Sun et al., 2023; Zeng et al., 2024; Zhang et al., 2024c), which discretize document embeddings into token sequences.
+Notable examples include numeric IDs (Tay et al., 2022; Wang et al., 2022; Zhuang et al., 2023a; Mehta et al., 2023) and codebook-derived tokens (Yang et al., 2023; Sun et al., 2023; Zeng et al., 2024b; Zhang et al., 2024c), which discretize document embeddings into token sequences.
 Existing generative retrieval methods within these two categories have continued to evolve, showing promising performance.
@@
 Leveraging this metric, we conduct extensive experiments to uncover the scaling behaviors of generative retrieval under different model sizes and data sizes.
-Additionally, we analyze how increased inference-time computation influences recall, providing further insight into the trade-offs and advantages of scaling computation.
+Additionally, we analyze how increased inference-time computation influences performance, highlighting its role in improving retrieval accuracy.
 From our extensive exploration, several intriguing findings stand out.
@@
 2) Expanding the training data scale benefits all methods, and n-gram-based approaches demonstrate especially robust gains, indicating a strong synergy between LLMs and natural identifiers.
-3) We find that boosting inference computation (e.g., FLOPs) yields clear performance improvements that follow power-law trends, revealing that generative retrieval can significantly profit from additional inference computation—an aspect rarely discussed in prior work.
-4) We discover that LLaMA models consistently outperform T5 models and exhibit higher theoretical upper bounds, hinting that the generative ability of larger decoder-only models may be particularly advantageous for generative retrieval.
+3) We discover that LLaMA models consistently outperform T5 models and exhibit higher theoretical upper bounds, hinting that the generative ability of larger decoder-only models may be particularly advantageous for generative retrieval.
+4) We find that boosting inference computation yields clear performance improvements that follow power-law trends, revealing that generative retrieval can significantly profit from additional inference computation—an aspect rarely discussed in prior work.
 2. Related Work
-In this section, we revisit the background of generative retrieval and scaling laws.
+In this section, we revisit the previous studies of generative retrieval and scaling laws.
 We first present the key advancements in generative retrieval, focusing on document identifier design and training methodologies.
@@
 2.1. Generative Retrieval
-Generative retrieval leverages generative models, particularly LLMs, reformulating the retrieval task as an auto-regressive generation problem.
-Document identifiers, as the core component, should be distinctive enough to represent related passages, reducing useless information in a passage and making it easier for the model to memorize and learn.
-The design of DocID can be broadly categorized into two approaches: natural identifiers and learned identifiers.
+Generative retrieval reformulates retrieval as an autoregressive generation task, where a language model directly generates the identifier of the target document given a query.
+A central component in generative retrieval is the document identifier (DocID), which serves as the generation target.
+DocID can be broadly categorized into two types: natural identifiers and learned identifiers.
 The first type uses identifiers that naturally carry semantic meaning about its associated document, which are comparably cost-efficient to be created, without the need for additional human supervision or forcing any structure in the search space.
@@
 Specifically, numeric ID (Tay et al., 2022; Wang et al., 2022; Zhuang et al., 2023a; Mehta et al., 2023) was found easy to construct but required extra memory steps.
-While codebook-derived tokens (Yang et al., 2023; Sun et al., 2023; Zeng et al., 2024; Zhang et al., 2024c) have been proven effective because the search space is reduced after each decoding step.
+And codebook-derived tokens (Yang et al., 2023; Sun et al., 2023; Zeng et al., 2024b; Zhang et al., 2024c) have been proven effective because the search space is reduced after each decoding step.
 Generative retrieval research has also advanced in training methods, which can be divided into generative training and discriminative training.
-In generative training, the model is trained to generate the appropriate identifier for a given query (Pradeep et al., 2023; Tay et al., 2022; Wang et al., 2022; Zeng et al., 2024; Bevilacqua et al., 2022), aligning naturally with the generative capabilities of LLMs to produce accurate document identifiers.
-Discriminative training, on the other hand, employs ranking losses (Li et al., 2024; Tang et al., 2024) and negative sample mining techniques (Zeng et al., 2024) to teach the model to produce a ranked list of documents.
-These approaches directly optimize retrieval performance by aligning the training objectives with the ranking requirements of retrieval tasks.
+In generative training, the model is trained to generate the appropriate identifier for a given query (Pradeep et al., 2023; Tay et al., 2022; Wang et al., 2022; Zeng et al., 2024b; Bevilacqua et al., 2022), aligning naturally with the generative capabilities of LLMs to produce accurate document identifiers.
+Discriminative training, on the other hand, employs ranking losses (Li et al., 2024b; Tang et al., 2024) and negative sample mining techniques (Zeng et al., 2024b) to teach the model to produce a ranked list of documents.
+These approaches align the training objectives with the ranking requirements of retrieval tasks.
 2.2. Neural Scaling Laws
@@
 OpenAI (Kaplan et al., 2020) extended this concept to larger models, demonstrating that model scaling yields consistent improvements in tasks like language modeling.
-Google (Hoffmann et al., 2022) further advanced this understanding by introducing a unified formula for scaling laws, thus laying the groundwork for scaling strategies in neural networks.
+Google (Hoffmann et al., 2022) further introduced a unified formula for scaling laws, thus laying the groundwork for scaling strategies in neural networks.
 Scaling laws have been successfully applied to various domain-specific fields such as speech recognition (Radford et al., 2023), computer vision (Dehghani et al., 2023; Zhai et al., 2022), and vision-language models (Jia et al., 2021; Radford et al., 2021).
-In the field of Information Retrieval, scaling laws have been explored in recommendation.
+In the field of information retrieval, scaling laws have been explored in recommendation (Lin et al., 2024).
 Studies have examined applications in Click-Through Rate (CTR) prediction (Ardalani et al., 2022) and sequential recommendation models (Zhang et al., 2024a) using unique item identifiers.
@@
 Another study on industrial multi-stage advertisement retrieval systems (Wang et al., 2024) emphasizes task-specific optimizations but lacks exploration of general scaling laws applicable to generative retrieval.
-In contrast, an early exploration of generative retrieval examines performance across varying corpus and model sizes (Pradeep et al., 2023) but is limited to learned identifier methods, leaving natural identifier methods unaddressed.
+An early exploration of generative retrieval examines performance across varying corpus and model sizes (Pradeep et al., 2023) but is limited to learned identifier methods, leaving natural identifier methods unaddressed.
 Additionally, it only adopts different sizes of T5 models and does not fit power-law scaling relationships.
-These gaps motivate our comprehensive study, which examines generative retrieval under varying model sizes, data scales, and DocID strategies, aiming to shed light on broader scaling behaviors and guide future system design.
-Following these discussions on generative retrieval and neural scaling laws, in this work, we conduct a comprehensive investigation of scaling behaviors in generative retrieval.
-By including larger decoder-only models and analyzing representative approaches, we aim to uncover key principles of scaling and provide insights to advance the design and optimization of generative retrieval systems.
+In contrast to prior work, our study investigates generative retrieval with larger decoder-only models, compares both learned and natural DocID strategies, and examines whether power-law relationships hold across model and data scales.
+Moreover, we analyze inference-time scaling behavior, which has received little attention in previous studies.
+These distinctions enable a more comprehensive understanding of generative retrieval scaling beyond existing task-specific or partial analyses.
 3. Methodology
 In this section, we first formalize the generative retrieval task in Section 3.1.
-Next, we describe the representative approaches for natural identifier and learned identifier: MINDER (Li et al., 2023b) and RIPOR (Zeng et al., 2024), respectively, in Section 3.2.
+Next, we describe the representative approaches for natural identifier and learned identifier: MINDER (Li et al., 2023b) and RIPOR (Zeng et al., 2024b), respectively, in Section 3.2.
 The backbones and training configurations used for generative retrieval are introduced in Section 3.3.
@@
 For natural identifiers, we use the n-gram-based method as n-grams effectively capture diverse semantic relationships between queries and documents.
-For learned identifiers, we adopt a codebook-based approach, as it leverages advanced neural methods to encode semantic information more compactly and effectively.
+For learned identifiers, we adopt a codebook-based approach, as it leverages advanced neural methods to encode semantic information effectively.
 3.2.1. N-gram-based Generative Retrieval
@@
 In this method, an LLM is trained to generate query-specific n-grams, which act as identifiers for ranking documents.
-The training process begins by selecting n-grams from documents based on their overlap or semantic similarity with user queries.
-These n-grams serve as the basis for training the LLM to predict relevant identifiers when presented with a query.
+As shown in Figure 2 (1), the training process begins by selecting n-grams from documents based on their overlap or semantic similarity with user queries.
+These n-grams serve as the basis for training the LLM to predict relevant identifiers given a query.
 At inference, the LLM generates n-grams for a query, and these are used to score and rank documents based on a heuristic function, such as n-gram frequency or semantic similarity.
@@
 Codebooks originate from techniques designed to create discrete visual data representations (Lee et al., 2022a; van den Oord et al., 2017).
-Learned codebooks for documents represent them as sequences of unique codes that effectively capture the semantics of their associated content (Yang et al., 2023; Sun et al., 2023; Zeng et al., 2024; Zhang et al., 2024c).
-The process begins by encoding documents into dense vector representations using an encoder network.
+Learned codebooks for documents represent them as sequences of unique codes that effectively capture the semantics of their associated content (Yang et al., 2023; Sun et al., 2023; Zeng et al., 2024b; Zhang et al., 2024c).
+As shown in Figure 2 (2), the process begins by encoding documents into dense vector representations using an encoder network.
 These dense vectors are then discretized into tokens by mapping them to entries in the learned codebook.
-Finally, a decoder network reconstructs the original document from the codebook tokens to ensure the generated representations are accurate and compact.
-The resulting code sequences serve as unique identifiers for documents, establishing a one-to-one correspondence between the sequence and the document.
+Finally, a decoder network reconstructs the original document from the codebook to ensure the generated representations are accurate and compact.
+The resulting code sequences serve as unique identifiers for documents, establishing a one-to-one correspondence between the code sequence and the document.
 At inference, the LLM generates a code sequence for a query, which is matched to the corresponding document based on the learned codebook.
-In our study, we select the RIPOR (Zeng et al., 2024) as the representative of the codebook-based method.
+In our study, we select the RIPOR (Zeng et al., 2024b) as the representative of the codebook-based method.
+Figure 2. Comparison of n-gram-based and codebook-based methods. N-grams are extracted text spans, while codebook-based methods are discrete representations generated via codebook.
 3.3. Backbone Models and Training Setting
@@
 3.3.1. Backbone Models
-For our experiments, we use the widely adopted T5 (Raffel et al., 2020) series as the primary backbone, which has been extensively employed in previous generative retrieval studies (Tay et al., 2022; Zeng et al., 2024; Ren et al., 2023; Mehta et al., 2023; Zhuang et al., 2023b; Zhang et al., 2024c).
+For our experiments, we use the widely adopted T5 (Raffel et al., 2020) series as the primary backbone, which has been extensively employed in previous generative retrieval studies (Tay et al., 2022; Zeng et al., 2024b; Ren et al., 2023; Mehta et al., 2023; Zhuang et al., 2023b; Zhang et al., 2024c).
 To evaluate the effect of model size, we experiment with all T5 variants: T5-Small, T5-Base, T5-Large, T5-XL, and T5-XXL, which differ only in parameter sizes while maintaining identical pre-training configurations.
@@
 This simplification is made because our study emphasizes scaling behavior rather than achieving absolute performance, and the body text serves as the most important identifier type.
-For each document, we select 10 n-grams, each consisting of 10 tokens, based on their overlap with the associated query to ensure high semantic relevance.
-The final training set consists of nearly 600,000 query-to-n-gram pairs.
+For each document, we select 10 n-grams, each consisting of 10 tokens, based on their overlap with the associated query to ensure semantic relevance.
+The final training set includes nearly 600,000 query-to-n-gram pairs.
 During training, the input consists of the query, and the label is a single n-gram from the corresponding document.
@@
 where $n$ represents an individual n-gram and $\theta$ represents the model parameters.
-Since each n-gram is treated independently, the model learns to predict each query-relevant n-gram as a separate target during training.
+Since each n-gram is treated independently, the model learns to predict each query-relevant n-gram as a separate target.
 To keep consistent with the following setup of codebook-based methods, we train each model for one epoch using the MINDER-provided data with LoRA (Hu et al., 2022).
 As recommended in MINDER, the learning rate for T5 models is set to 3e-5.
-For LLaMA models, we use a learning rate of 3e-4, a commonly adopted value for fine-tuning LLaMA.
+For LLaMA models, we use a learning rate of 3e-4, a commonly adopted value.
 $\bullet$ Codebook-Based Generative Retrieval.
-Following the selected representative codebook-based generative retrieval method, RIPOR (Zeng et al., 2024), we conduct experiments using the MSMARCO-1M dataset, a subset of the MSMARCO (Nguyen et al., 2017) dataset, containing one million passages and query-document pairs.
+Following the selected representative codebook-based generative retrieval method, RIPOR (Zeng et al., 2024b), we conduct experiments using the MSMARCO-1M dataset, a subset of the MSMARCO (Nguyen et al., 2017) dataset, containing one million passages and query-document pairs.
 This dataset is chosen because RIPOR is only available on MSMARCO, and since our study focuses on scaling behavior rather than absolute performance, the choice of the dataset has minimal impact on the relative results.
@@
 Building on the concept of contrastive entropy used in dense retrieval, we adapt it to the generative setting by incorporating the loss associated with query-to-identifier generation.
+$\bullet$ Contrastive generation loss.
 In generative retrieval, the primary objective is to generate identifiers that correspond to relevant documents while distinguishing them from irrelevant ones.
@@
 where $\mathcal{L}(q,\mathcal{I})$ represents the generation loss for the query $q$ to produce the identifier $\mathcal{I}$, calculated using the cross-entropy loss.
-The contrastive generation loss is a continuous and highly sensitive metric designed to effectively evaluate retrieval performance across different generative retrieval methods and models.
-1) For n-gram-based generative retrieval, where each document is represented by multiple identifiers, the $\mathcal{L}(q,\mathcal{I}_{d^{+}})$ is extended by averaging the generation loss across all n-grams associated with the positive document.
-2) For codebook-based generative retrieval, the query-to-document relationship is evaluated through the loss of generating the document’s unique code sequence.
-3) Furthermore, this metric is versatile across different LLM architectures, as it directly leverages the generative capabilities of the models to compute losses.
-4) Moreover, the CGL is based on the relative proportion of the positive document’s loss compared to negative documents, reducing sensitivity to the absolute magnitude of the loss and ensuring consistency across varying model scales and training settings.
+Unlike contrastive loss (Khosla et al., 2020) or contrastive entropy (Fang et al., 2024) which operates in embedding space, CGL directly uses generation loss to assess the model’s preference for positive identifiers.
+It requires no additional training objective and works purely in the generation setting.
+CGL has the following advantages:
+  • Method compatibility.
+CGL supports both n-gram-based and codebook-based generative retrieval.
+For n-gram identifiers, it averages the generation loss overall n-grams of the positive document.
+For codebook methods, it evaluates the loss of generating the complete token sequence assigned to each document.
+  • Model compatibility. CGL is applicable across different LLMs, as it operates solely on forward generation loss without requiring model-specific structures or additional training objectives.
+  • Relative evaluation. By using the ratio between positive and negative losses, CGL mitigates sensitivity to absolute loss values, ensuring consistent evaluation.
+$\bullet$ Validation.
 To validate the proposed CGL, we conducted experiments to analyze its relationship with existing ranking metrics, such as Recall, NDCG, MAP, and MRR.
@@
 4. Training Scaling Laws
-In this section, we present the results of our experiments and summarize our initial investigation into the scaling laws for generative retrieval.
-Specifically, we analyze how model size, training data size, and identifier methods influence retrieval performance, using contrastive generation loss as the evaluation metric.
-These experiments aim to uncover the relationships and trends that define the scalability of generative retrieval systems.
+In this section, we present the results of our experiments and summarize our investigation into the training scaling laws for generative retrieval.
+Specifically, we analyze how model size, training data size, and identifier methods influence retrieval performance, using CGL as the evaluation metric.
 4.1. Model Size Scaling
-We now investigate the impact of model size on generative retrieval performance. We begin by examining n-gram-based generative retrieval.
+We now investigate the impact of model size on retrieval performance. We begin by examining n-gram-based generative retrieval.
 4.1.1. N-Gram-Based Generative Retrieval
+We investigated the effect of model size on generative retrieval performance using the n-gram-based method.
+Models of varying sizes, including T5 and LLaMA, were fine-tuned on query-to-n-gram training pairs, and their performance was evaluated using the CGL on the test set.
+Figure 4 (1) illustrates the scaling behavior of T5 and LLaMA models concerning this metric.
+Based on the observed relationship between model size and the CGL, we propose a scaling law to quantify this behavior as follows:
+(6) $\mathcal{L}_{\text{CGL}}(P)=\left(\frac{\gamma}{P}\right)^{\alpha}+\lambda_{P}.$
+Here, $P$ represents the number of non-embedding parameters of the model, and $\mathcal{L}_{\text{CGL}}(P)$ denotes the CGL on the test set.
+The parameters $\gamma$, $\alpha$, and $\lambda_{P}$ are coefficients determined through fitting.
+Here, $\lambda_{P}$ represents the irreducible loss, a theoretical lower bound on performance as $P$ approaches infinity, accounting for limitations such as dataset noise and variability in relevance judgments.
+Using the least squares method, we fit the scaling law and report the fitted coefficients for T5 and LLaMA in Table 1.
+The results reveal several important insights: 1) Both models demonstrate a strong power-law relationship between model size and contrastive generation loss, with exceptionally high coefficients of determination.
+2) LLaMA demonstrates comprehensive performance advantages over T5, characterized by a more efficient scaling mechanism.
+Specifically, LLaMA achieves lower CGL across model sizes and exhibits a steeper improvement curve (scaling exponent of $\alpha=2.40$ versus T5’s $\alpha=0.40$).
+LLaMA’s smaller irreducible loss ($\lambda_{P}$) suggests a higher potential performance ceiling, indicating its superior capability to approach the theoretical limits of generative retrieval performance.
+These findings highlight LLaMA’s promising performance and signal the potential of decoder-only architectures.
 Table 1. Fitted parameters for the scaling law on model sizes with n-gram-based methods.
@@
 N-Gram LLaMA Series $1.24\times 10^{8}$ $2.40$ $0.00328$ $0.999$
-We investigated the effect of model size on generative retrieval performance using the n-gram-based method.
-Models of varying sizes, including T5 and LLaMA, were fine-tuned on query-to-n-gram training pairs, and their performance was evaluated using the contrastive generation loss on the test set.
-Figure 3 (1) illustrates the scaling behavior of T5 and LLaMA models concerning this metric.
-Based on the observed relationship between model size and the contrastive generation loss, we propose a scaling law to quantify this behavior as follows:
-(6) $\mathcal{L}_{\text{CGL}}(P)=\left(\frac{\gamma}{P}\right)^{\alpha}+\lambda_{P}.$
-Here, $P$ represents the number of non-embedding parameters of the model, and $\mathcal{L}_{\text{CGL}}(P)$ denotes the contrastive generation loss on the test set.
-The parameters $\gamma$, $\alpha$, and $\lambda_{P}$ are coefficients determined through fitting.
-Here, $\lambda_{P}$ represents the irreducible loss, a theoretical lower bound on performance as $P$ approaches infinity, accounting for limitations such as dataset noise and variability in relevance judgments.
-Using the least squares method, we fit the scaling law and report the fitted coefficients for T5 and LLaMA in Table 1.
-The results reveal several important insights: 1) Both models demonstrate a strong power-law relationship between model size and contrastive generation loss, with exceptionally high coefficients of determination (T5: R² = 0.996, LLaMA: R² = 0.999).
-2) LLaMA demonstrates comprehensive performance advantages over T5, characterized by a more efficient scaling mechanism.
-Specifically, LLaMA achieves lower contrastive generation loss across model sizes and exhibits a steeper improvement curve (scaling exponent of $\alpha=2.40$ versus T5’s $\alpha=0.40$).
-Notably, LLaMA’s smaller irreducible loss ($\lambda_{P}$) suggests a higher potential performance ceiling, indicating its superior capability to approach the theoretical limits of generative retrieval performance.
-These findings highlight LLaMA’s promising performance and potentially signal the broader potential of decoder-only architectures.
 4.1.2. Codebook-Based Generative Retrieval
@@
 We leave this possibility for future research as the computational intensity needed makes comprehensive exploration impractical for us.
-Despite the lack of scaling trends, LLaMA consistently achieves lower contrastive generation loss than T5 across all model sizes, highlighting its stronger retrieval capabilities.
+Despite the lack of scaling trends, LLaMA consistently achieves lower CGL than T5 across all model sizes, highlighting its stronger retrieval capabilities.
 These findings suggest LLaMA demonstrates promising performance characteristics and may signal avenues for future research in generative retrieval.
@@
 The results presented in Figure 4 highlight clear differences between the n-gram-based and codebook-based methods in terms of scaling behavior and overall performance.
-1) The n-gram-based method significantly outperforms the codebook-based approach, demonstrating substantially lower contrastive generation loss.
+1) The n-gram-based method significantly outperforms the codebook-based approach, demonstrating lower CGL.
 Even the most advanced LLaMA models utilizing codebook tokens cannot match the performance of T5 models with n-gram-based retrieval.
@@
 In this section, we investigate how varying the training data size influences retrieval performance while keeping other factors, such as model size and architecture, constant.
-Table 2. Fitted parameters for the scaling law on data sizes.
-Method Model $\eta$ $\beta$ $\lambda_{D}$ $R^{2}$
-N-Gram LLaMA-7B $1.05\times 10^{4}$ $3.99$ $0.00335$ $0.990$
-Codebook LLaMA-7B $9.28\times 10^{-1}$ $1.01$ $0.00389$ $0.983$
 4.2.1. N-Gram-Based Generative Retrieval
@@
 (7) $\mathcal{L}_{\text{CGL}}(D)=\left(\frac{\eta}{D}\right)^{\beta}+\lambda_{D}$
-Here, $D$ represents the number of query-identifier pairs, and $\mathcal{L}_{\text{CGL}}(D)$ denotes the contrastive generation loss on the test set.
+Here, $D$ represents the number of query-identifier pairs, and $\mathcal{L}_{\text{CGL}}(D)$ denotes the CGL on the test set.
 The parameters $\eta$, $\beta$, and $\lambda_{D}$ are coefficients determined through fitting.
@@
 Using the least squares method, we fit the scaling law to the observed data, achieving a coefficient of determination of $R^{2}=0.990$, which indicates a strong fit.
-As seen in Figure 4 (1), retrieval performance improves significantly as the training data size increases, with the contrastive generation loss decreasing sharply.
+As seen in Figure 5 (1), retrieval performance improves significantly as the training data size increases, with the CGL decreasing sharply.
 The power-law scaling behavior reflects the model’s capacity to leverage larger datasets to better capture the semantic relationships between queries and identifiers.
+Table 2. Fitted parameters for the scaling law on data sizes.
+Method Model $\eta$ $\beta$ $\lambda_{D}$ $R^{2}$
+N-Gram LLaMA-7B $1.05\times 10^{4}$ $3.99$ $0.00335$ $0.990$
+Codebook LLaMA-7B $9.28\times 10^{-1}$ $1.01$ $0.00389$ $0.983$
 4.2.2. Codebook-Based Generative Retrieval
@@
 The fitted curve in Figure 5 (2) achieves a coefficient of determination of $R^{2}=0.983$, indicating a strong fit.
-As the training data size increases, the contrastive generation loss decreases steadily, demonstrating that retrieval performance improves with larger datasets.
-The results highlight that even for the codebook-based method, which involves learning entirely new representations unrelated to the model’s pretraining objectives, increasing the data size leads to significant retrieval performance enhancements.
+As the training data size increases, the CGL decreases steadily, demonstrating that retrieval performance improves with larger datasets.
+The results highlight that even for the codebook-based method, which involves learning entirely new representations unrelated to the model’s pretraining objectives, increasing the data size leads to performance enhancements.
 4.2.3. Comparsions.
@@
 The low scaling exponent implies that this method requires substantially more training data to achieve comparable performance.
-Recent studies suggest more advanced training strategies, such as ranking losses (Tang et al., 2024; Zeng et al., 2024), could potentially address these learning challenges.
+Recent studies suggest more advanced training strategies, such as ranking losses (Tang et al., 2024; Zeng et al., 2024b), could potentially address these learning challenges.
 4.3. Model-Data Joint Laws
@@
 5. Inference Scaling Laws
-In the previous sections, we demonstrated the existence of scaling laws in generative retrieval concerning model size and data size, highlighting how larger models and richer data yield stronger performance.
-These findings align with trends observed in LLMs, further validating the synergy between generative retrieval and LLMs.
-Recent studies have also revealed another dimension of scaling: the scaling of computational investment during inference (Wu et al., 2024).
-Specifically, increasing inference computing, such as scaling the number of decoding tokens, has been shown to yield substantial performance gains.
-This trend suggests that inference scaling may be a critical factor in optimizing LLM-based systems, complementing improvements from larger models and data.
-In generative retrieval, inference scaling is particularly promising because the identifiers used for retrieval are generated dynamically at this stage.
-Generative retrieval inherently involves generating query-specific identifiers to score and rank documents.
-The quality and diversity of generated identifiers are directly influenced by the inference process, including parameters such as beam size, which controls the breadth of the search during generation.
-Larger beam sizes allow for the exploration of a broader search space, potentially improving retrieval performance, but they also increase inference latency, creating a trade-off between quality and efficiency.
-Motivated by this, we investigate whether the inference scaling laws observed in LLMs extend to generative retrieval.
-Given that generative retrieval shares a strong alignment with the LLMs, we wonder whether similar scaling behaviors will hold for generative retrieval inference.
+In the previous sections, we demonstrated the existence of training scaling laws in generative retrieval concerning model size and data size.
+Beyond training, recent studies have also revealed another dimension of scaling: the scaling of computational investment during inference (Wu et al., 2024).
+Specifically, increasing inference computing, such as scaling the number of decoding tokens, yielded substantial performance gains.
+This raises the question of whether similar benefits extend to generative retrieval.
+Inference scaling is particularly promising in generative retrieval because the identifiers used for retrieval are generated dynamically at this stage.
+The generation process, controlled by parameters such as beam size, directly affects the quality and diversity of identifiers, and hence retrieval performance.
+Motivated by these observations, we investigate whether inference scaling laws observed in LLMs also hold in generative retrieval, and how factors like beam size and compute budget impact its effectiveness.
 Figure 6.
@@
 This increase not only boosts the quantity of n-grams available for scoring but also enhances their diversity and quality, which is likely to improve the final document retrieval performance.
+$\bullet$ Implementation details.
 We use the n-gram-based generative retrieval method described in Section 3.2.1.
@@
 During inference, we vary the beam size, testing values of $B=\{1,5,10,20,50,100\}$ and record the corresponding inference computation needed.
-Beam size determines the number and search space of candidate n-grams generated per query, which in turn affects the diversity of identifiers used for document retrieval.
+Beam size determines the number and search space of candidate n-grams generated per query, affecting the diversity of identifiers used for document retrieval.
 Increasing the beam size effectively increases the computational cost during inference, as a larger search space requires more floating-point operations (FLOPs) to generate candidate n-grams.
-To ensure comparability, the number of retrieved documents is fixed at 100 for all configurations, and other settings like batch size remain constant.
+$\bullet$ Evaluation.
 To evaluate retrieval performance, we define a metric called Miss Rate (MR), which measures the proportion of relevant documents that are not retrieved within the top $k$ results.
@@
 5.2. Results
-The results of our experiments are summarized in Figure 5, which illustrates the Miss Rate across different inference FLOPs per query for both T5-Base and LLaMA-7B models.
+The results of our experiments are summarized in Figure 6, which illustrates the MR across different inference FLOPs per query for both T5-Base and LLaMA-7B models.
 We evaluate MR at different retrieval thresholds ($k=5,20,100$) to assess the retrieval performance under varying levels of precision.
@@
 where $C$ represents the inference FLOPs per query, $\mu$, $\sigma$, and $\lambda_{C}$ are parameters to fit, and $\lambda_{C}$ is irreducible loss, a theoretical lower bound on retrieval performance as $C$ approaches infinity.
-The fitted curves in Figure 5, reveal a consistent trend for both T5-Base and LLaMA-7B models: the Miss Rate decreases steadily as the inference FLOPs per query increase.
+The fitted curves in Figure 6, reveal a consistent trend for both T5-Base and LLaMA-7B models: the MR decreases steadily as the inference FLOPs per query increase.
 This validates the effectiveness of inference scaling in improving generative retrieval performance.
-The decrease in MR is particularly pronounced at lower computational budgets, where increased beam sizes lead to significant performance gains.
-As shown in Table 3, the proposed scaling law provides an excellent fit to the experimental data for both models, with high $R^{2}$ values in most cases.
-Additionally, the rate of decline in MR varies with the retrieval threshold $k$.
-For both models, MR@5 decreases the fastest, reflecting the substantial benefits of inference scaling for high-precision retrieval.
-Conversely, MR@100 exhibits a slower decline, indicating that broader recall settings are less sensitive to increased computational resources.
-Comparing the two models, when inference FLOPs are below $10^{11}$, T5-Base outperforms LLaMA-7B, achieving lower Miss Rates across all $k$ values.
-However, as FLOPs increase beyond $10^{11}$, LLaMA-7B demonstrates a clear advantage, with all three Miss Rates achieving lower values.
-This highlights that LLaMA-7B gains substantially more performance rise from increased inference computation compared to T5-Base, making it highly effective in high-computation scenarios.
-Notably, LLaMA-7B’s irreducible loss ($\lambda_{C}$) for MR@5 is significantly lower than T5-Base, indicating superior potential for high-precision retrieval, while its irreducible loss for MR@20 and MR@100 is slightly higher, suggesting T5-Base retains an advantage in broader recall settings.
-These results highlight the trade-offs between the two models, with LLaMA-7B excelling in high-precision tasks, achieving lower Miss Rates for the top-ranked results as computational budgets grow, while T5-Base demonstrates greater potential for broader recall tasks like larger-scale retrieval.
+As shown in Table 3, the proposed scaling law fits well with the experimental results.
+The results reveal several key findings:
+  • MR@k sensitivity.
+The rate of MR decline varies with the retrieval threshold $k$.
+MR@5 decreases the fastest for both models, showing strong benefits from inference scaling in high-precision retrieval, while MR@100 declines more slowly, indicating reduced sensitivity in broader recall settings.
+  • Model performance under low compute.
+When inference FLOPs are below $10^{11}$, T5-Base achieves lower MR across all $k$ than LLaMA-7B, suggesting that T5-Base is more efficient in compute-constrained scenarios.
+  • Model performance under high compute.
+As FLOPs increase beyond $10^{11}$, LLaMA-7B gradually surpasses T5-Base, achieving lower MR@100 and MR@20, demonstrating greater benefits from increased inference computation.
+  • Irreducible loss.
+LLaMA-7B exhibits lower irreducible loss ($\lambda_{C}$) for MR@5, suggesting superior capacity for high-precision retrieval.
+In contrast, it shows a slightly higher irreducible loss for MR@20 and MR@100, indicating that the T5-Base retains an advantage in broader recall scenarios.
 6. Discussion
-In this work, we explored the training scaling laws of model size, data size, and inference scaling laws in generative retrieval, proposing metrics to analyze retrieval performance under varying computational budgets.
-By systematically varying these factors, we highlighted how scaling influences retrieval performance and identified key trade-offs between inference computation and performance.
-While our findings provide valuable insights into generative retrieval systems, several important aspects remain unexplored and warrant further investigation.
-We introduced a new metric contrastive generation loss to evaluate the generative retrieval capability of models.
-While CGL provides a relative measure of a model’s retrieval ability under the same settings, it is not equivalent to traditional metrics such as Recall.
-For example, two models with similar CGL values may exhibit differing recall performance, and a lower CGL does not guarantee higher recall.
-CGL is useful for comparing models in relative terms within the same experimental setup, but its interpretability as an absolute measure of retrieval performance is limited.
-Another limitation of this work lies in the training methodology, as we relied solely on the basic generative cross-entropy loss.
-While this aligns well with n-gram-based methods, which match the model’s pretraining objective and require minimal fine-tuning, it poses challenges for codebook-based methods that need to learn entirely new tokens or relationships, demanding more training data and epochs.
-We attempted to use RIPOR’s full dataset, containing over 80 million query-to-code sequences from MSMARCO, but this scale of data was computationally infeasible for LLaMA, as it would require thousands of hours for us to complete the training.
-Future research with greater resources could further explore the potential of such extensive datasets and more advanced training objectives.
-In our experiments, we observed that codebook-based methods did not exhibit a clear scaling effect with increasing model size, but showed a pronounced scaling effect with training data size.
-This discrepancy may stem from the significantly higher learning difficulty of codebook representations compared to n-grams.
-The amount of training data used in our experiments may have been insufficient, leaving the models in an underfitted state and preventing the scaling effect with model size from emerging.
-Previous studies have shown that model performance can undergo substantial improvement at a certain point (Du et al., 2024), it is possible that with sufficient training data and time, codebook-based methods might also exhibit scaling with model size.
-Future research with enhanced settings and resources could further investigate this possibility.
-In this study, we primarily focused on the n-gram-based method for inference scaling.
-We observed improved retrieval performance by increasing beam size and validated its alignment with power-law scaling.
-However, codebook-based methods could also potentially benefit from inference scaling strategies, such as increasing beam size to explore a larger search space.
-Due to space limitations, we did not explore this direction, but it represents a possible direction for future research.
+In this work, we explored the training and inference scaling laws in generative retrieval, proposing metrics to analyze retrieval performance under varying conditions.
+While our findings offer useful insights, several limitations and open questions remain.
+$\bullet$ Limitation of CGL.
+While the proposed Contrastive Generation Loss reflects the model preference for correct identifiers under consistent settings, it does not directly correlate with traditional metrics like Recall.
+For example, models with similar CGL values may have different recall, and a lower CGL does not guarantee better performance.
+CGL is effective for relative comparisons, but its interpretability as an absolute performance measure is limited.
+$\bullet$ Simplified training settings.
+For simplicity, we adopt the standard generative cross-entropy loss without incorporating more advanced objectives.
+While this may be sufficient for n-gram-based methods, which match the model’s pretraining objective, it poses challenges for codebook-based methods that need to learn entirely new tokens or relationships, demanding more training data and epochs.
+Prior work has explored more effective alternatives, such as learning-to-rank (LTR) losses (Li et al., 2024b; Tang et al., 2024), which may better capture the retrieval objective.
+Future research could revisit this direction with more advanced training objectives.
+$\bullet$ Model scaling of the codebook-based method.
+In our experiments, codebook-based methods showed clear scaling with training data size but not with model size.
+This may be due to the greater learning difficulty of codebook representations.
+The training data and time in our experiments may have been insufficient, preventing the scaling effect from emerging.
+Previous studies have shown that model performance can undergo substantial improvement at a certain point (Du et al., 2024).
+Although we attempted to use RIPOR’s full dataset (over 80M query-code pairs), the computational cost for LLaMA was unaffordable.
+Future work could revisit this setting to better understand the scaling potential of codebook-based methods.
+$\bullet$ Inference scaling of the codebook-based method.
+Our inference scaling experiments focused on n-gram-based retrieval, showing improved performance with increased beam size and alignment with power-law scaling.
+Codebook methods may also benefit from inference-time strategies such as beam search (Zeng et al., 2024a), but we leave this exploration to future work due to space and resource constraints.
+$\bullet$ Dataset inconsistency across methods.
+In our experiments, the n-gram-based and codebook-based methods were trained and evaluated on different datasets, as we directly adopted identifier sets from prior work to simplify implementation and isolate scaling effects.
+Consequently, the comparison between methods may be affected by dataset differences.
+A more controlled comparison would involve evaluating both methods on a unified dataset—a direction we leave for future work.
 7. Conclusion and Future Work
-In this paper, we explored the scaling laws of generative retrieval, analyzing how model size, training data size, model architecture, and retrieval methods impact retrieval performance.
-Using Contrastive Generation Loss as our primary metric, we found a clear power-law relationship between model size and retrieval performance for n-gram-based methods.
+In this paper, we explored the scaling laws of generative retrieval across training and inference dimensions.
+We analyzed how model size, training data size, and retrieval methods influence performance.
+Using Contrastive Generation Loss as the metric, we found a clear power-law relationship between model size and retrieval performance for n-gram-based methods.
 Additionally, we observed consistent scaling trends with increasing data size for both n-gram-based and codebook-based methods.
 A comparison between architectures showed that LLaMA consistently outperformed T5 under identical experimental conditions, highlighting LLaMA’s superior capability for advancing generative retrieval.
-Beyond model and data scaling, we also investigated inference scaling, specifically for n-gram-based methods.
-Our results revealed that increasing inference computation followed a distinct power-law trend, significantly boosting retrieval performance.
-This unique inference scaling offers complementary opportunities for optimizing retrieval efficiency and performance.
-Collectively, our findings validate the effectiveness of systematically scaling model capacity, training data, and inference computation, suggesting promising pathways to further enhance generative retrieval methods.
-While our research provides a foundation for understanding the scaling laws of generative retrieval, several directions for future exploration remain.
-First, further investigation into scaling laws for codebook-based methods is warranted, particularly in the context of model size.
-Exploring advanced training objectives could better capture the potential scaling benefits of these methods, which were not fully realized under our experimental settings.
-Second, with access to greater computational resources, revisiting extensive datasets could enable deeper investigations into the scaling laws of generative retrieval, particularly for methods that require larger-scale training.
-Finally, while this study focused on n-gram-based methods for inference scaling, exploring similar scaling strategies for codebook-based methods, such as increasing beam size, remains an intriguing direction for future research.
+Beyond training scaling, we also investigated inference scaling for n-gram methods and found that increasing inference-time computation followed a power-law trend, offering a complementary axis for performance gains.
+Collectively, our findings validate the effectiveness of systematically scaling model capacity, training data, and inference computation, suggesting pathways to further enhance generative retrieval methods.
+While our study offers a foundation for understanding the scaling laws of generative retrieval, several directions remain for future exploration.
+First, adopting more effective training objectives, such as learning-to-rank losses, may improve retrieval performance by better aligning model optimization with the retrieval task.
+Second, for codebook-based methods, the limited model-size scaling observed in our experiments suggests a need for larger training datasets and longer training schedules to fully realize their potential.
+Finally, although our inference scaling analysis focused on n-gram-based methods, codebook-based methods may also benefit from increased inference-time computation, such as using larger beam sizes—offering a valuable direction for extending the scope of inference scaling in generative retrieval.
+Acknowledgements.
+This research is supported by the National Research Foundation, Singapore under its National Large Language Models Funding Initiative (AISG Award No: AISG-NMLP-2024-002), Research Grants Council of Hong Kong(PolyU/15209724) and PolyU internal grants (BDWP).
+Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not reflect the views of National Research Foundation, Singapore.